OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point shown and is perpendicular to the line shown.

OpenStudy (imstuck):

The y intercept is where the line goes through the y axis so that is -6, and the slope is up 3 and over 1, so the equation in y = mx+b form is y = 3x-6. The line shown has an equation of y = 3x-6. The slope of the new line that will go through the point is the negative reciprocal of that, making it -1/3x. Now we just need to use the point given to find the equation.

OpenStudy (imstuck):

\[y-y _{1}=m(x-x _{1})\]where m is your slope of -1/3 and the x1 and y1 coordinates are from your point which is ((6,-6). So fill in accordingly:

OpenStudy (imstuck):

\[y-(-6)=-1/3x(x-6)\]Expanding that you will get:

OpenStudy (imstuck):

\[y+6=-\frac{ 1 }{ 3 }x+2\]and \[y=-\frac{ 1 }{ 3 }x+2-6\]so\[y=-\frac{ 1 }{ 3 }x-4\]

OpenStudy (imstuck):

Do you see the steps that get you where you want to be?

OpenStudy (anonymous):

yes that makes so much more sense now thanks for the help I understand in way more.@IMStuck

OpenStudy (anonymous):

y = ax + b => if 0 = a * 2 + b if x = 0 => b = -6 so : y = 3x -6 's 90 degree is -1/3 => y = -(1/3) x +b if corss the point (6,-6) = > -6 = -(1/3) * 6 + b => b = -4 so : y = -(1/3) x -4